# Flavor diagonal nucleon charges using clover fermions on MILC HISQ ensembles

### v2

@@ VI.1 Axial
CCFV fits in Fig. 8 and 9 for $g_{A}^{u}$ and $g_{A}^{d}$, respectively, show similar large positive slope versus $M_{\pi}^{2}$, but with a larger uncertainty in the “$N\pi$” analysis. There also is a significant positive slope in $g_{A}^{u}$ versus $a$ but not in $g_{A}^{d}$. These variations decrease $g_{A}^{u}$ and make $g_{A}^{d}$ more negative as $a\to 0$ and $M_{\pi}\to 135$ MeV. They add in $g_{A}^{u+d}$ and almost cancel in $g_{A}^{u-d}$.
@@
Estimates of $g_{A}^{d}$ are consistent between the two renormalization methods but $\approx 10\%$ more negative with the “$N\pi$” analysis. Based on the reasoning given in the extraction of the isovector axial form factors in [11, 12], the “$N\pi$” analysis is suggested, however, given the size of the errors, we defer choosing. So we take the mean and quote half the difference, $0.024$, as a systematic uncertainty along with $0.009$ for the difference between the the CCFV and CC values. This gives $g_{A}^{d}=-0.440(29)(9)_{\rm CC}(24)_{N\pi}$.
@@ VI.2 Tensor
The magnitude of ESC in $g_{T}^{u}$ is significant, while the errors in $g_{T}^{d}$ and $g_{T}^{s}$ are large, especially on the physical pion mass ensemble $a09m130$, as shown in Fig. 2 and in Ref. [3]. The CCFV fits are shown in Fig. 11, 12 and 13. The “standard” and “$N\pi$” analyses give consistent estimates but the errors in the “$N\pi$” case are larger. Since no enhanced contributions from multihadron, “$N\pi$”, states is expected, we choose the “standard” analysis and the CCFV fits. For all three charges, $g_{T}^{u,d,s}$, we average the values from the two renormalization methods and use the difference between CCFV and CC fits to assign a systematic uncertainty. The final results are summarized in Tables 7 and 8.
@@ VI.3 Scalar
The behavior of $g_{S}^{u}$ and $g_{S}^{d}$ versus $a$ is similar and there is a significant negative slope, i.e., the value increases as $a\to 0$ as shown in Figs. 14 and 15. Similarly, the value increases as $M_{\pi}\to 135$ MeV. There is a striking difference in the chiral behavior between the “standard” and “$N\pi$” fit values, and the extrapolation in the $N\pi$ data is large. The data point $a09m130$ at the physical pion mass, therefore, plays a crucial role in the extrapolation. Results after CCFV extrapolation of the “standard” and “$N\pi$” analyses data are given in Tables 6 and 7. Based on the $\chi$PT analysis presented in Ref. [4], we pick the “$N\pi$” analysis for the final values of $g_{S}^{u,d}$ given in Table 8. Again, we choose the “standard” analysis value for $g_{S}^{s}$ as the lowest multihadron state is expected to be $\Sigma K$.
@@ VI.4 The pion-nucleon sigma term $\sigma_{\pi N}$
Results for the renormalization group invariant pion-nucleon sigma term were presented in Ref. [4] where it was calculated on each ensemble using the bare quantities $\sigma_{\pi N}=m_{l}^{\text{bare}}g_{S}^{u+d,\text{bare}}$ and the data were extrapolated to the physical point using the N2LO $\chi$PT expression [39]:
@@
Our $\chi$PT analysis suggested that all five terms contribute significantly in a CC fit [4]. With data at only three values of $M_{\pi}\approx 135,220,310$ MeV, we can only explore the chiral part of the ansatz with a maximum of three terms. We make a second fit by including one more term, $\propto d_{3}^{\chi}$, by fixing its coefficient to the $\chi$PT value evaluated with $M_{N}=0.939\mathop{\rm GeV}\nolimits$, $g_{A}=1.276$, and $F_{\pi}=92.3\mathop{\rm MeV}\nolimits$ [4]. That paper also showed that the neglected finite-volume corrections can also be estimated in $\chi$PT, indicating a correction of less than $1\mathop{\rm MeV}\nolimits$ for the physical pion mass ensemble $a09m130$. The data and fits are shown in Fig. 17. The updated results, $\sigma_{\pi N}|_{N\pi}=61(6)$ MeV for the “$N\pi$” analysis and $\sigma_{\pi N}|_{N\pi}=41(6)$ MeV for the “standard” analysis, overlap with those given in Ref. [4].
In this work, we also calculate $\sigma_{\pi N}$ and $\sigma_{s}$ using ESC fits to the sum and do the CCFV extrapolation using the renormalized scalar charges $g_{S}^{u+d}$ and $g_{S}^{u,d,s}$ [^1] Note that using the individual values for $g_{S}^{u,d}$ from Table 7 gives a $\sigma_{\pi N}$ that is about 2 MeV larger. to obtain the renormalized isoscalar charge $g_{S}^{u+d}=12.4(1.2)$ and $g_{S}^{s}=0.37(14)$ for the “standard” analysis and $g_{S}^{u+d}=17.8(1.6)$ for the “$N\pi$” analysis. Combining these with the renormalized quark masses $m_{l}=3.427(51)$ and $m_{s}=93.46(58)$ MeV for the $N_{f}=2+1+1$ theory given in the FLAG 2024 report [5], our results are: $\sigma_{\pi N}|_{N\pi}=61(6)$ MeV for the $N\pi$ analysis, and $\sigma_{\pi N}|_{\rm standard}=43(4)$ MeV and $\sigma_{s}=35(13)$ MeV from the “standard” analysis.
Averaging these two estimates for $\sigma_{\pi N}$ and taking the larger of the errors, our final values are
$\sigma_{\pi N}|_{\rm standard}$ $=42(6)~{}{\rm MeV}$
$\sigma_{\pi N}|_{N\pi}$ $=61(6)~{}{\rm MeV}$
@@
Extrapolation including finite volume corrections (CC fit). $\rm{Z_{1}}$ method
@@
Extrapolation including finite volume corrections (CC fit). $\rm{Z_{2}}$ method
@@
Table 6: Results for the flavor diagonal charges obtained after extrapolation to the physical point with and without finite volume corrections (the CCFV and CC fits). Results are presented for the two strategies (“standard” and “$N\pi$”) for removing ESC and the two methods ($\rm{Z_{1}}$ and $\rm{Z_{2}}$, defined in Appendix B.4) for determining the renormalization constants. Note the tiny difference in $g_{T}^{s}$ between the “standard” and “$N\pi$” analysis since the mixing from light flavors depends on $N\pi$.
@@
$d$ $-0.440(29)(9)_{\rm CC}(24)_{N\pi}$ $-0.195(10)(2)_{\rm CC}$ $8.84(93)(1)_{\rm CC}$ $6.04(63)(1)_{\rm CC}$ $-0.438(18)(30)$ $-0.204(11)(10)$
$s$ $-0.055(9)(1)_{\rm CC}$ $-0.0016(12)(1)_{\rm CC}$ $0.66(17)(5)_{\rm CC}$ $0.37(13)(6)_{\rm CC}$ $-0.053(8)$ $-0.027(16)$
Table 7: Final results for the flavor diagonal charges obtained by combining the values given in Table 6 using criteria specified in the text. The subscript CC denotes the systematic uncertainty assigned using the difference between the CCFV and CC fit values; and $N\pi$ for that between the the “standard” and “$N\pi$” strategies where appropriate. Values with the $\rm{Z_{1}}$ and $\rm{Z_{2}}$ renormalization methods (defined in Appendix B.4) are averaged. Results for $g_{A}^{q}$ and $g_{T}^{q}$ supersede those published previously in Refs. [2] and [3] and reproduced here.
@@
$d$ $-0.440(39)$ $-0.195(10)$ $8.84(93)$
$s$ $-0.055(9)$ $-0.0016(12)$ $0.37(14)$
Table 8: Final results for the flavor diagonal charges obtained by combining, in quadrature, the errors given in Table 7. The results for the axial charges, $g_{A}^{u,d}$, are the average over the “$N\pi$” and “standard” analysis; the tensor, $g_{T}^{u,d}$, and the strange charges, $g_{A,S,T}^{s}$, are from the “standard” analysis; and $g_{S}^{u,d}$ are from the “$N\pi$” analysis as explained in the text.
@@ VI.5 Extracting the isovector charges $g_{A,T,S}^{u-d}$
The disconnected contributions, which have large errors, cancel exactly in the isovector charges, $g_{A,T,S}^{u-d}$, for the isospin-symmetric theory. Thus, our overall strategy to get the isovector charges is to make ES fits to the difference of the connected contributions, i.e., to $C^{\text{u,conn}}_{\Gamma}(t;\tau)-C^{\text{d,conn}}_{\Gamma}(t;\tau)$. These fits are much more robust compared to those for the flavor diagonal charges, and take into account correlated fluctuations that are missed in the construction $g^{u,bare}-g^{d,bare}$. An example of fits to $C^{\text{conn+disc}}_{\Gamma}(t;\tau)$ giving $O(1)$ errors is the scalar charge $g_{S}^{u-d}$. This occurs for the two reasons illustrated by the $a09m130$ data in Fig. 5. First, the data for the larger $\tau$ ($\tau=14$ and $16$) do not satisfy the two conditions that must hold for fits to give consistent results—monotonic convergence from below and symmetry about $\tau/2$. These features are exhibited by the data from the other seven ensembles and we have checked that they improve with increasing statistical precision. Second, within errors, which increase rapdly with $\tau$, the data around $\tau/2$ overlap for $\tau=8,10,12$. This, ideally, indicates convergence by the plateau criteria, however, for confirmation much higher precision $\tau=14$ and $16$ data are needed. On the other hand, using 2-state fits to remove residual ESC not well-constrained.
@@
Our final results for isovector charges from the full set of clover-on-HISQ calculations (using 13 data points on 11 HISQ ensembles, CCFV extrapolation, and taking the average of $\rm{Z_{1}}$ and $\rm{Z_{2}}$ values) have already been published in Ref. [12]. For completeness, Table 9 summarizes the values from this work using eight ensembles, “standard” and “$N\pi$” analyses, CCFV and CC extrapolations; and the same two renormalization methods. The two sets of results are consistent, with the values obtained here having slightly larger errors. The only notable feature is the difference between the various estimates for $g_{A}^{u-d}$ that has already been pointed out in Ref. [12].
@@
Table 9: Results for the isovector charges with the two strategies, “standard” and “$N\pi$”, used to remove ESC; the two methods, $\rm{Z_{1}}$ and $\rm{Z_{2}}$, used for renormalization; and the two extrapolation ansatz, CCFV and CC, to the physical point. The final values from this work are compared with the average of $\rm{Z_{1}}$ and $\rm{Z_{2}}$ values from Ref. [12] obtained with CCFV extrapolation of the full set of thirteen clover-on-HISQ calculations. The CCFV extrapolations based on eight versus thirteen points account for most of the differences between the results. The “standard” analysis strategy is labeled “$3^{\ast}$” in Ref. [12], while the “$N\pi$” analysis is similar to “3-RD”.
Figure 6: Comparison with other lattice calculations using 2+1+1 or 2+1-flavors of quarks that meet the FLAG criteria. Unfilled symbols indicate that these results have not yet been published or are conference proceedings. The references are ETM 24 [40] PNDME 18A [2] PNDME 18B [3] Mainz 19A [41] $\chi$QCD 18 [42]
Figure 7: (Left) Results for $\sigma_{\pi N}={m}_{ud}g_{S}^{u+d}$ from 2+1- and 2+1+1-flavor lattice calculations that satisfy the FLAG criteria for inclusion in averages. The BMW 20 1+1+1+1-flavor result is still a preprint and listed here under 2+1+1-flavor calculations for brevity. We follow the FLAG conventions and denote direct determinations by squares and the results obtained using the Feynman-Helmann method by triangles. The references for the lattice results are: ETM 24 [40], Gupta 21ahb [4], BMW 20A [43] ETM 14A [44], Mainz 23ehe [45] RQCD 22xux [46] $\chi$QCD 15A [47], BMW 15 [48], BMW 11A [43]. The phenomenological estimates (blue filled circles), obtained using $\pi N$ scattering data, are from Hoferichter 23 [49], Ruiz de Elvira 17 [50], Hoferichter 15 [51], Chen 12 [52] Alarcon 11 [53]. (Right) Results for $\sigma_{s}$. The additional references for $\sigma_{s}$ are MILC 12C [54], Junnarkar 13 [55], MILC 09D [56]
@@ VII Comparison with Previous Results and Conclusions
From the axial charges, $g_{A}^{u,d,s}$ given in Table 8, we calculate the contribution of the intrinsic spin of quarks to the nucleon spins to be $0.5\Delta\Sigma_{q}=0.5\sum_{q}g_{A}^{q}=0.143(24)$. This is consistent with the COMPASS analysis $0.180.13<0.5\Delta\Sigma_{q}<$ 0.18 evaluated at 3 GeV2 [57].
@@
Our results for the scalar charges, $g_{S}^{u,d,s}$, are new with only preliminary versions presented in conference proceedings [6, 7]. Most of the discussion of scalar charges in literature has been in terms of the sigma terms, the pion-nucleon, $\sigma_{\pi N}$, and the strangeness content $\sigma_{s}$ [61]. For the pion-nucleon sigma term $\sigma_{\pi N}$, the main issue, first raised in Ref. [4], is removing the contributions of multihadron $N\pi$ and $N\pi\pi$ states that are lattice artifacts. Our selection of the “$N\pi$” analysis results, as discussed in Section VI.4, rely on the $\chi$PT analysis and the phenomenological observation of a large scalar condensate in QCD. Results in Eq. (10) are consistent with those presented in Ref. [4]. The value $\sigma_{s}=35(13)$ MeV reflects the large uncertainty in the calculation of $g_{S}^{s}$ and the CCFV extrapolation as shown in Fig. 16. A comparison with other lattice and phenomenological results in the FLAG format is shown in Fig. 7. Our result $\sigma_{\pi N}|_{\rm standard}=42(6)~{}{\rm MeV}$ is consistent with other lattice measurements, while our preferred value $\sigma_{\pi N}|_{N\pi}=61(6)~{}{\rm MeV}$ is in agreement with phenomenological extraction (dispersive analysis of $\pi N$ scattering data).
Note that a larger value of the scalar charges, output of our “$N\pi$” analyses motivated by $\chi$PT, has a significant impact on the reach of dark matter direct-detection experiments—a $\approx 50\%$ larger $g_{S}^{u,d}$ increases the cross-section by a factor of $\approx 2.25$ assuming the couplings to dark matter is in the most promising scalar channel. In fact, all three sets of charges, $g_{A,S,T}^{u,d,s}$, give the coupling of dark matter or novel Higgs-like interactions to nucleons in the respective Lorentz channels.
@@
  • Our data show ESC in $g_{A}^{l,s}$ (see Figs. 2 and 3) while the ETM data (labelled $g_{A}^{u+d(disc)}$ and $g_{A}^{s(disc)}$) are flat even though they include data at smaller $\tau$, i.e., in the range $1.60.41<\tau<$ 1.6 fm versus $1.50.9<\tau<$ 1.5 fm in our calculation. This accounts for most of the difference in the final values of $g_{A}^{u,d}$ shown in Fig. 6.
  • Both calculations find large ESC in $g_{S}^{u,d,s}$. The difference in the estimates of $\sigma_{\pi N}$ comes from the choice of ESC fits. We have chosen the “$N\pi$” fit for the final answer based on our $\chi$PT analysis [4]. Had we chosen our “standard” analysis value, as ETM does, the two results would be consistent.
@@
Looking ahead, we are analyzing two physical pion mass HISQ ensembles at $a\approx 0.087$ and $0.057$ fm with about six times larger statistics (6000 and 4000 configurations, respectively) and better tuned values of $m_{s}$ and $m_{l}$ in the clover action. These calculations will improve control over excited-state analysis and, hopefully, the statistics are large enough to provide a data driven resolution between the “standard” and “$N\pi$” analyses.

### v3

@@ VI.1 Axial
CCFV fits in Fig. 8 and 9 for $g_{A}^{u}$ and $g_{A}^{d}$, respectively, show similar large positive slope versus $M_{\pi}^{2}$, but with a larger uncertainty in the $N\pi$” analysis. There also is a significant positive slope in $g_{A}^{u}$ versus $a$ but not in $g_{A}^{d}$. These variations decrease $g_{A}^{u}$ and make $g_{A}^{d}$ more negative as $a\to 0$ and $M_{\pi}\to 135$ MeV. They add in $g_{A}^{u+d}$, making it smaller, and almost cancel in $g_{A}^{u-d}$.
@@
Estimates of $g_{A}^{d}$ are consistent between the two renormalization methods but $\approx 10\%$ more negative with the $N\pi$” analysis. Based on the reasoning given in the extraction of the isovector axial form factors in [11, 12], the $N\pi$” analysis is suggested, however, given the size of the errors, we defer choosing. So we take the mean and quote half the difference, $0.024$, as a systematic uncertainty along with $0.009$ for the difference between the the CCFV and CC values. This gives $g_{A}^{d}=-0.440(29)(9)_{\rm CC}(24)_{N\pi}$.
@@ VI.2 Tensor
The magnitude of ESC in $g_{T}^{u}$ is significant, while the errors in $g_{T}^{d}$ and $g_{T}^{s}$ are large, especially on the physical pion mass ensemble $a09m130$, as shown in Fig. 2 and in Ref. [3]. The CCFV fits are shown in Fig. 11, 12 and 13. The “standard” and $N\pi$” analyses give consistent estimates but the errors in the $N\pi$” case are $\approx 50\%$ larger. For all three charges, $g_{T}^{u,d,s}$, we average the values from the two renormalization methods and use the difference between CCFV and CC fits to assign a systematic uncertainty. Since no enhanced contributions from multihadron, $N\pi$”, states is expected in the tensor channel, our choice is the “standard” analysis and the CCFV fits. Nevertheless, to be conservative and account for the $\approx 2\sigma$ difference between the “standard” and $N\pi$ values for $g_{T}^{d}$, we assign half the difference, 0.012, as an additional systematic uncertainty. The final results are summarized in Tables 7 and 8.
@@ VI.3 Scalar
[…]
@@ VI.4 The pion-nucleon sigma term $\sigma_{\pi N}$
[…]
@@
[…]
In this work, we also do the CCFV extrapolation of the renormalized charges $g_{S}^{u+d}$ and $g_{S}^{u,d,s}$. Our preferred value for the isoscalar charge, $g_{S}^{u+d}$, is obtained by doing ESC fits to the sum of the total $u$ and $d$ correlators. The results are $g_{S}^{u+d}=12.4(1.2)$ for the “standard” and $g_{S}^{u+d}=17.8(1.6)$ for the $N\pi$” analysis. The value $g_{S}^{s}=0.37(14)$ is, as explained before, taken from the standard analysis. Combining these with the renormalized quark masses $m_{l}=3.427(51)$ and $m_{s}=93.46(58)$ MeV given in the FLAG 2024 report [5] for the $N_{f}=2+1+1$ theory, provides the second estimates: $\sigma_{\pi N}|_{N\pi}=61(6)$ MeV for the $N\pi$ analysis, and $\sigma_{\pi N}|_{\rm standard}=43(4)$ MeV and $\sigma_{s}=35(13)$ MeV from the “standard” analysis. Note that if the individual values of $g_{S}^{u}$ and $g_{S}^{d}$ are taken from Table 7, then $\sigma_{\pi N}$ is larger by about 2 MeV.
Averaging the two estimates for $\sigma_{\pi N}$ and taking the larger of the errors, our final values are
$\sigma_{\pi N}|_{\rm standard}$ $=42(6)~{\rm MeV}\,$,
$\sigma_{\pi N}|_{N\pi}$ $=61(6)~{\rm MeV}\,$,
@@
Extrapolation excluding finite volume corrections (CC fit). $\rm{Z_{1}}$ method
@@
Extrapolation excluding finite volume corrections (CC fit). $\rm{Z_{2}}$ method
@@
Table 6: Results for the flavor diagonal charges obtained after extrapolation to the physical point with and without finite volume corrections (the CCFV and CC fits). Results are presented for the two strategies (“standard” and $N\pi$”) for removing ESC and the two methods ($\rm{Z_{1}}$ and $\rm{Z_{2}}$, defined in Appendix B.4) for determining the renormalization constants. Note the tiny difference in $g_{T}^{s}$ between the “standard” and $N\pi$” analysis since the mixing from light flavors depends on $N\pi$.
@@
$d$ $0.440(29)(9)_{\rm CC}(24)_{N\pi}$ $0.195(10)(2)_{\rm CC}(12)_{N\pi}$ $8.84(93)(1)_{\rm CC}$ $6.04(63)(1)_{\rm CC}$ 0.438(18)(30)-0.438(18)(30) 0.204(11)(10)-0.204(11)(10)
$s$ $055(9)(1)_{\rm CC} 0.0$$0016(12)(1)_{\rm CC}$ $0.66(17)(5)_{\rm CC}$ $37(13)(6)_{\rm CC} 0.0$53(8)-0.053(8) 0.027(16)-0.027(16)
Table 7: Final results for the flavor diagonal charges obtained by combining the values given in Table 6 using criteria specified in the text. The subscript $C$ denotes the systematic uncertainty assigned using the difference between the CCFV and CC fit values; and $N\pi$ for that between the “standard” and $N\pi$” strategies where appropriate. Values with the $\rm{Z_{1}}$ and $\rm{Z_{2}}$ renormalization methods (defined in Appendix B.4) are averaged. Results for $g_{A}^{q}$ and $g_{T}^{q}$ supersede those published previously in Refs. [2] and [3] and reproduced here.
@@
$d$ 0.440(39)-0.440(39) 0.195(16)-0.195(16) $8.84(93)$
$s$ 0.055(9)-0.055(9) 0.0016(12)-0.0016(12) $0.37(14)$
Table 8: Final results for the flavor diagonal charges obtained by combining, in quadrature, the errors given in Table 7. The results for the axial charges, $g_{A}^{u,d}$, are the average over the $N\pi$” and “standard” analysis; the tensor, $g_{T}^{u,d}$, and the strange charges, $g_{A,S,T}^{s}$, are from the “standard” analysis; and $g_{S}^{u,d}$ are from the $N\pi$” analysis as explained in the text.
@@ VI.5 Extracting the isovector charges $g_{A,T,S}^{u-d}$
The disconnected contributions, which have large errors, cancel exactly in the isovector charges, $g_{A,T,S}^{u-d}$, for the isospin-symmetric theory. Thus, our overall strategy to get the isovector charges is to make ES fits to the difference of the connected contributions, i.e., to $C^{\text{u,conn}}_{\Gamma}(t;\tau)-C^{\text{d,conn}}_{\Gamma}(t;\tau)$. These fits are much more robust compared to those for the individual flavor diagonal charges, and benefit from correlated fluctuations that are missed in the construction $g^{u,bare}-g^{d,bare}$. For example, extracting $g_{S}^{u-d}$ from separate fits to the two $C^{\text{conn+disc}}_{S}(t;\tau)$ gives $O(1)$ errors with current data. This occurs for the two reasons illustrated by the $a09m130$ data in Fig. 5. First, the data for the larger $\tau$ ($\tau=14$ and $16$) do not satisfy the two conditions that must hold for fits to give consistent results—monotonic convergence from below and symmetry about $\tau/2$. These features are exhibited by the data from the other seven ensembles and we have checked that they improve with increasing statistical precision. Second, within errors, which increase rapidly with $\tau$, the data around $\tau/2$ overlap for $\tau=8,10,12$. This, ideally, indicates convergence by the plateau criteria, however, for confirmation much higher precision $\tau=14$ and $16$ data are needed. On the other hand, using 2-state fits to remove residual ESC not well-constrained.
@@
Our final results for isovector charges from the full set of clover-on-HISQ calculations (using 13 data points on 11 HISQ ensembles, CCFV extrapolation, and taking the average of values with $\rm{Z_{1}}$ and $\rm{Z_{2}}$ renormalization) have already been published in Ref. [12]. For completeness, Table 9 summarizes the values from this work using eight ensembles, “standard” and $N\pi$” analyses, CCFV and CC extrapolations; and the same two renormalization methods. The two sets of results are consistent, with the values obtained here having slightly larger errors. The only notable feature in both is the difference between the various estimates for $g_{A}^{u-d}$ that has already been pointed out in Ref. [12].
@@
Table 9: Results for the isovector charges with the two strategies, “standard” and $N\pi$”, used to remove ESC; the two methods, $\rm{Z_{1}}$ and $\rm{Z_{2}}$, used for renormalization; and the two ansatz, CCFV and $C$, used to extrapolate data to the physical point. The final values from this work are compared with the average of $\rm{Z_{1}}$ and $\rm{Z_{2}}$ values from Ref. [12] obtained with CCFV extrapolation of the full set of thirteen clover-on-HISQ calculations. The CCFV extrapolations based on eight versus thirteen points account for most of the differences between the results. The “standard” analysis strategy is labeled $3^{\ast}$” in Ref. [12], while the $N\pi$” analysis is similar to “3-RD”.
[…]
@@ VII Comparison with Previous Results and Conclusions
From the axial charges, $g_{A}^{u,d,s}$ given in Table 8, we calculate the contribution of the intrinsic spin of quarks to the nucleon spins to be $0.5\Delta\Sigma_{q}=0.5\sum_{q}g_{A}^{q}=0.143(24)$. This is consistent with the COMPASS analysis $0.13<0.5\Delta\Sigma_{q}<0.18$ evaluated at 3 GeV2 [57].
@@
[…]
@@
  • Our data show ESC in $g_{A}^{l,s}$ (see Figs. 2 and 3) while the ETM data (labelled $g_{A}^{u+d(disc)}$ and $g_{A}^{s(disc)}$) are flat even though they include data at smaller $\tau$, i.e., in the range $0.41<\tau<1.6$ fm versus $0.9<\tau<1.5$ fm in our calculation. This accounts for most of the difference in the final values of $g_{A}^{u,d}$ shown in Fig. 6.
[…]
@@
Looking ahead, we are analyzing two physical pion mass HISQ ensembles at $a\approx 0.087$ and $0.057$ fm with about six times larger statistics (6000 and 4000 configurations, respectively) and better tuned values of $m_{s}$ and $m_{l}$ in the clover action. These calculations will improve control over excited-state analysis and, hopefully, the statistics are large enough to provide a data driven resolution between the “standard” and $N\pi$” analyses.